OpenStudy (anonymous):

Find the distance from the point (-3,2,1) to the plane 1x+5y-2z=-3

OpenStudy (anonymous):

find a set of parametric equations for the line through the point that is perpendicular to the plane, then find the intersection of the line and the plane and plug this into the parametric equation

OpenStudy (anonymous):

You want to find any vector that goes from the point to the plane. In this case it is <-3-x, 2-y, 1-z>. Since the normal vector is be the shortest distance, you want to project onto that normal vector. The normal vector is <1, 5, -2> in this case. The magnitude of that projection is the shorted distance from the plane to the point. Do you know how to project a vector?

OpenStudy (anonymous):

Yeah I got the formula around here somewhere

OpenStudy (anonymous):

wio can you help me out, because im not really sure what im doing

OpenStudy (anonymous):

@math_proof with what?

OpenStudy (anonymous):

with that proof

OpenStudy (anonymous):

(u*v/||u^2||)*u

OpenStudy (anonymous):

that i posted

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://mathinsight.org/distance_point_plane

OpenStudy (anonymous):

@ConDawg, still need help?

OpenStudy (anonymous):

Yea, so if Normal vector is <1,5,-2>, and wwe want to project it on <-3,2,1>?

OpenStudy (anonymous):

We want to project <-3-x, 2-y, 1-z> onto it.

OpenStudy (anonymous):

And we want to find the magnitude of that projection.

OpenStudy (anonymous):

So we take the dot product of <-3-x, 2-y, 1-z> and <1,5,-2> first of all.

OpenStudy (anonymous):

Then we divide that by |<1,5,-2>|, the magnitude of the normal vector.

OpenStudy (anonymous):

So 5/sqrt(30)

OpenStudy (anonymous):

No, the dot product should have x, y, and z in it Where did they go?

OpenStudy (anonymous):

i was taught to always give a new points not the magnitude of it

OpenStudy (anonymous):

We'll get rid of them soon enough, but unless you're ahead of the game, they should still be in there.

OpenStudy (anonymous):

@ConDawg What is the dot product of <-3-x, 2-y, 1-z> and <1,5,-2>?

OpenStudy (anonymous):

The dot product gives a scaler, which is just a single number, is what my text book says, so should I ignore that part for now?

OpenStudy (anonymous):

We will get a scalar, it's just that it should have x, y, and z in it.

OpenStudy (anonymous):

wio can you give me some guide to do part b of my proof?

OpenStudy (anonymous):

I get 1(-3-x) + 5(2-y) - 2(1-x)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Now when you expand that you get?

OpenStudy (anonymous):

-3-x+10-5y-2+2x

OpenStudy (anonymous):

Okay, all the like terms need to be added up...

OpenStudy (anonymous):

Like -3, 10, and -2 are like terms.

OpenStudy (anonymous):

-x-5y+2x+5

OpenStudy (anonymous):

Now we use the fact that 1x+5y-2z=-3 to get rid of our x, y and z.

OpenStudy (anonymous):

If 1x+5y-2z=-3 then -x-5y+2z = 3

OpenStudy (anonymous):

Okay got cha

OpenStudy (anonymous):

So it's \(\Large \frac{3+5}{\sqrt{30}} or \frac{8\sqrt{30}}{30} \)

OpenStudy (anonymous):

I hate this stuff

OpenStudy (anonymous):

LOL, me too!

OpenStudy (anonymous):

Nice its right thanks a bunch!!!!